please help me solve my homework: simplify: sin(90degrees+ x) sin ( 180 degrees +x) + cos ( 90 degrees + x) cos ( 180 degrees - x ).
Answer by stanbon(48564)

simplify: sin(90degrees+ x) sin ( 180 degrees +x) + cos ( 90 degrees + x) cos ( 180 degrees - x ).
------------
There are several ways to work this problem; you might want to do it anothe way,
but here goes:
sin(90+x) = cosx
sin(180+x) = -sinx
cos(90+x) = -sinx
cos(180-x) = -cosx
--------------------
Your Problem:
cosx*(-sinx)+(-sinx)(-cosx)
==0
============
Cheers,
Stan H.